OpenStudy (anonymous):

Does anyone know how to solve this? csc^4(2X)-4=0 for all real numbers?

mathslover (mathslover):

Hello tolli, first of all Welcome to OpenStudy.

mathslover (mathslover):

First of all let me know that what you tried yet to solve this question ?

OpenStudy (anonymous):

so, first I wrote csc in terms of 1/sin. so this is what I first wrote:1/sin^4x(2x)-4. Then I added the 4 to the other side: 1/sin^4(2x)=4. Then when I tried to factored/take out a greatest common factor in 1/sin^4x but it just got messy I I got lost.

mathslover (mathslover):

OK! wait let me check your method

mathslover (mathslover):

what I will prefer you to do is : \[\large{\frac{1}{\sin^4(2x)} = 4}\] \[\large{1 = 4\sin ^4 (2x)}\] \[\large{\sqrt{1}= \sqrt{4\sin^4(2x)}}\] What do you get now ?

OpenStudy (anonymous):

+/-1=+/-2sin^2(2x)

mathslover (mathslover):

OK leave \(\pm\) for once. Now ?

OpenStudy (anonymous):

would I divide by 2 then square everything?

mathslover (mathslover):

OK see here : \[\large{1 = 2 \sin^2(2x)}\] \[\large{\frac{1}{2} = \sin^2(2x)}\] Right ?

mathslover (mathslover):

Now, square root ...

mathslover (mathslover):

Not square but square root

OpenStudy (anonymous):

YES. then you square it to get: 1/4=sin(2x)

OpenStudy (anonymous):

oh ok

mathslover (mathslover):

square root it and then tell me what you get

OpenStudy (anonymous):

sqrt2/2=sin(2x)

mathslover (mathslover):

Correct, well done!. Or I can write that as : \[\large{\frac{1}{\sqrt{2}} = \sin(2x)}\]

OpenStudy (anonymous):

should I multiply sqrt2/2 by 2?

mathslover (mathslover):

No. Not now

mathslover (mathslover):

Can you tell me at which angle \(\large{\sin (x) }\) is \(\large{\frac{1}{\sqrt{2}}}\) ?

OpenStudy (anonymous):

45

OpenStudy (anonymous):

and 2pi/4

mathslover (mathslover):

45 degrees is right but \(\large{\frac{2\pi}{4}}\) is not as 2pi/4 = pi/2 which is equal to 90 degrees not 45 degrees friend. So you have 45 degree as pi/4 radians. ok?

OpenStudy (anonymous):

ok

mathslover (mathslover):

Now, we have : \[\large{\frac{1}{\sqrt{2}} = \sin(2x)}\] We know that : \[\large{\sin (\frac{\pi}{4}) = \frac{1}{\sqrt{2}}}\] right ?

mathslover (mathslover):

so can you tell me what will be "x" now ?

OpenStudy (anonymous):

pi/2

mathslover (mathslover):

how ?

OpenStudy (anonymous):

ok so, sin(pi/4)=1/srqt2, so if I double that I get 2pi/4 which simplifys to pi/2.

mathslover (mathslover):

oh! You have a misconception there . see : \[\large{\sin(\frac{\pi}{4}) = \frac{1}{\sqrt{2}}}\] also \[\large{\sin (2x) = \frac{1}{\sqrt{2}}}\] therefore : \[\large{\sin(\frac{\pi}{4}) = \sin(2x)}\] as they both are equal to the same thing therefore pi/4 = 2x

mathslover (mathslover):

can you tell me now , what is x equal to ?

OpenStudy (anonymous):

can I take sin inverse of pi/4 then divide by 2?

mathslover (mathslover):

See you can do like this : \[\large{\sin(\frac{\pi}{4}) = \frac{1}{\sqrt{2}}}\] \[\large{\frac{\pi}{4} = \sin^{-1}(\frac{1}{\sqrt{2}})}\] now divide by 2

OpenStudy (anonymous):

what happen to the sin(2x)?

mathslover (mathslover):

now : \[\large{\sin(2x) = \frac{\pi}{4}}\] \[\large{2x = \sin^{-1}(\frac{\pi}{4})}\] now divide by 2 . can you do that now?

OpenStudy (anonymous):

yes.but what is the sin inverse (pi/4)? Is it still pi/4?

mathslover (mathslover):

oh sorry mistake there : \[\large{\sin(2x) = \frac{1}{\sqrt{2}}}\] \[\large{2x = \sin^{-1}(\frac{1}{\sqrt{2}})}\] now divide by 2 , m sorry

OpenStudy (anonymous):

so, x= 2/sqrt2=2sqrt2/2=sqrt2

mathslover (mathslover):

no see now : \[\large{\sin^{-1}(\frac{1}{\sqrt{2}}) = \frac{\pi}{4}}\] therefore I have 2x = pi/4 x = pi/8

OpenStudy (anonymous):

oh. so when writhing the equation because the original question asked for all real numbers, will it be:x=pi/8+2npi

mathslover (mathslover):

npi +/- pi/8

OpenStudy (anonymous):

why is npi instead of 2npi?

mathslover (mathslover):

the formula for sin(pi/8) general soln is : \[\large{n\pi + (-1)^n (\frac{\pi}{8})}\]

OpenStudy (anonymous):

where is the -1^n coming from?

mathslover (mathslover):

http://mathsfirst.massey.ac.nz/Trig/TrigGenSol.htm

terenzreignz (terenzreignz):

I'm a bit disillusioned, here... Might I suggest an alternative way to do it? Starting from \[\huge \sin^2(2x)=\frac{1}{2}\]

OpenStudy (anonymous):

i would like that but I have to go to school. unless you can do it in 5min.

terenzreignz (terenzreignz):

Okay, let's see :) Then \[\huge \sin(2x)=\pm\frac{1}{\sqrt{2}}\]

OpenStudy (anonymous):

mathslovers thanks for all your help.

terenzreignz (terenzreignz):

What are all the angles from 0 to 2pi that have sines of plus or minus 1/sqrt(2) They are \[\frac{\pi}{4},\frac{3\pi}{4},\frac{5\pi}{4},\frac{7\pi}{4}\]

OpenStudy (anonymous):

yes. I understand that

terenzreignz (terenzreignz):

And these are possible values for 2x, so possible values for x are halves of these... \[\large \frac{\pi}{8},\frac{3\pi}{8},\frac{5\pi}{8},\frac{7\pi}{8}\] plus half of 2pi, which is pi

OpenStudy (anonymous):

so what is the general solution?

terenzreignz (terenzreignz):

These differ by pi/4, so let's add pi to pi/8, we get 9pi/8, which is just 7pi/8 + pi/4 Hence, if you add any multiple of pi/4 to pi/8, you still get a solution... The general solution is \[\huge \frac{\pi}{8}+\frac{k\pi}{4} \ \ \ \ k \in\mathbb{Z}\]

terenzreignz (terenzreignz):

That was a really quick summary, but that gets you there :)

OpenStudy (anonymous):

Thankyou so much your method help a lot.